OpenStudy (nightmarenight):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)?

OpenStudy (the_fizicx99):

Well bae, M = Y2-Y1/X2-X1 M = -5-(-1)/-2-3 M = -5 + 1/-5 M = -4/-5 Your slope is 4/5

OpenStudy (the_fizicx99):

*cough* medal *cough*

OpenStudy (smored):

slope=\[\frac{y _{2}-y _{1}}{x _{2}-x _{1}}\] \[\frac{ -5+1 }{ -2-3 }\] -4/-5

OpenStudy (the_fizicx99):

a negative and a negative = a ?? @smored

OpenStudy (smored):

-5-(-1)=-5+1

OpenStudy (smored):

yeah -4/-5=4/5

OpenStudy (nightmarenight):

Oh thank you <3 @tHe_FiZiCx99

OpenStudy (the_fizicx99):

:** muaa

OpenStudy (nightmarenight):

:****** yay @tHe_FiZiCx99